# Adversarial Attacks of Vision Tasks in the Past 10 Years: A Survey

### v1

@@ 7.5. Adversarial Defenses in LVLM
[(Li et al., 2024d; Zhang et al., 2024a; Chen et al., 2024b) , color=brightlavender, fill=brightlavender!15, text width=1.7cm, text=black]
@@
[(Ji et al., 2024b; Zhang et al., 2024a) , color=brightlavender, fill=brightlavender!15, text width=1.7cm, text=black]
@@
[(Zhang et al., 2023c; Li et al., 2024b) , color=brightlavender, fill=brightlavender!15, text width=1.7cm, text=black]
@@
[(Wu et al., 2023; Qi et al., 2024; Wang et al., 2024c; Gou et al., 2024) , color=brightlavender, fill=brightlavender!15, text width=1.7cm, text=black]
@@
[(Chen et al., 2023b; Phute et al., 2023; OpenAI, 2023c) , color=brightlavender, fill=brightlavender!15, text width=1.7cm, text=black]
@@
Fig. 9 illustrates the classification of adversarial defense methods in LVLM. Defense methods during the training phase involve fine-tuning the model or inputs to enhance robustness, while those during the inference phase focus on data detection and modification. As mentioned in JBA (Qi et al., 2024) and SSA-CWA (Dong et al., 2023a), FFT (Full Fine-Tuning) through adversarial training is computationally expensive. Therefore, the cost-effective methods like fine-tuning partial parameters with LoRA (Ji et al., 2024b; Zhang et al., 2024a), prompt tuning of the inputs (Zhang et al., 2023c; Li et al., 2024b), and the inference-phase strategies are more prevalent.
@@ 7.5.1. Training Phase
Inference-phase defense can be achieved by fine-tuning models or inputs. RTVLM (Li et al., 2024d), SPA-VL (Zhang et al., 2024a), and Dress (Chen et al., 2024b) apply adversarial FFT on LVLMs with SFT (Supervised Fine-Tuning), RLHF, and NLF-based CLF (Natural Language Feedback based Conditional Reinforcement Learning), respectively. To reduce computational costs, AdvLoRA (Ji et al., 2024b) and SPA-VL (Zhang et al., 2024a) explore LoRA-based adversarial training strategies. Inspired by prompt tuning, AdvPT (Zhang et al., 2023c) and APT (Li et al., 2024b) attempt to fine-tune the context of labels in image captioning under CLIP to improve robustness.
@@ 7.5.2. Inference Phase
Inference-phase methods can be further divided into pre-processing and post-processing, based on the time of LVLM’s replying. In pre-processing, JailGuard (Zhang et al., 2023d) detects whether the input is harmful by observing the LVLM’s response to different input variations (e.g., word changes or image transformations). ESCO (Gou et al., 2024) concatenates the LVLM’s self-generated image description to the original prompt to form a new input, avoiding adversarial information in the image. SASP (Wu et al., 2023) and AdaShield (Wang et al., 2024c) insert elaborate defense prompts into user prompts, guiding the model to generate safe content. JBA (Qi et al., 2024) finds that diffusion models (Nie et al., 2022) can effectively purify jailbreak images. In post-processing, LSD (Phute et al., 2023), CM-4 (OpenAI, 2023c), and PPI (Chen et al., 2023b) use additional LLMs or the victim moddel itself to detect whether the response is harmful.
@@ 8.1.1. Traditional Adversarial Attacks
In traditional adversarial attacks, transfer rates of untargeted attacks have reached around 90% (even with defenses) (Chen et al., 2023c), while targeted transfer rates lag significantly (Qin et al., 2022). Targeted attacks are more demanding, and improving their transferability is challenging. For physical robustness, attacks mainly focuses on target function design and transformation aggregation ways (section 5.2), which seems overly monotonous. Since physically robust AEs threaten many applications (section 5.2), discovering new ways to enhance this robustness offers valuable insights for defense. Moreover, if these samples also transfer across models, they pose an even greater threat to real-world systems. Physically robust samples often require visible perturbations, while increasing naturalness is a hot spot for improving stealthiness. These two may seem contradictory. Is there a way to combine both—creating seemingly natural but physically threatening samples without relying on camouflage? In addition, exploring potential adversarial threats in popular applications such as style transfer (Shan et al., 2023) and face-swapping (Dong et al., 2023b) is of practical significance.
@@ 8.1.2. Adversarial Attacks on LVLM
In adversariality, attacks on text modalities are less common than visual ones. Strengthening text-based attacks and exploring the vulnerability in cross-modality links could be promising directions. For LVLMs, both the visual encoder and LLM use Transformer architectures, potentially aiding transfer attacks, but cross-model transferability remains a challenge. As LVLMs are integrated into applications like autonomous driving, safety concerns may grow. Research on the physically robust AEs will likely extend into LVLM contexts. The huge scales and cost of commercial APIs make generating effective AEs in LVLMs more expensive, driving interest in cost-effective methods. LVLM downstream applications, like AI assistants (Wu et al., 2024) and robots, also face potential adversarial threats. Furthermore, while numerous safety related benchmarks (Shen et al., 2023; Liu et al., 2023d; Li et al., 2024a; Luo et al., 2024b; Zhang et al., 2024b; Li et al., 2024d; Tu et al., 2023; Sun et al., 2023b) for LVLMs have been proposed recently, there is still a lack of a widely accepted, unified, and comprehensive evaluation system.
@@ 8.2. Conclusion
This article reviews the development of visual adversarial attacks over the past decade. As Qi et al. note, adversarial attacks are shifting from classification-focused approaches to broader applications across LLMs. The article is thus divided into two parts: 1) traditional adversarial attacks, and 2) LVLM adversarial attacks.
The first part summarizes adversariality, transferability, and generalization, explaining the causes of adversariality and transferability, the roles of AEs, traits of transferability, and types of generalization. It then defines the problem and introduces threat models, victim models, datasets, and evaluation metrics, classifying traditional attacks into two phases: basic strategies, which explore various attack paradigms, and attack enhancement, aimed at boosting effectiveness. Motivations of the second phase fall into four types: improving transferability, physical robustness, stealthiness, and generation speed. This part concludes with an overview of the application of attacks across different tasks.
@@
Finally, the article discusses future research directions, including adversariality, transferability, physical robustness, stealth, generation speed, and applications, aiming to provide insights for future works in visual adversarial attacks.
@@ 9. Acknowledgments
We thank the Collaborative Innovation Center of Novel Software Technology and Industrialization for their support.
@@
- 3) (2017) NIPS 2017 Defense Competition (Rank 3). 2017. NIPS-r3. (2017). https://github.com/anlthms/nips-2017/tree/master/mmd
- Abid et al. (2021) Abubakar Abid, Maheen Farooqi, and James Zou. 2021. Large language models associate Muslims with violence. Nature Machine Intelligence (2021).
- Agrawal et al. (2019) Harsh Agrawal, Karan Desai, Yufei Wang, Xinlei Chen, Rishabh Jain, Mark Johnson, Dhruv Batra, Devi Parikh, Stefan Lee, and Peter Anderson. 2019. Nocaps: Novel object captioning at scale. In ICCV.
- Ahmed et al. (1974) Nasir Ahmed, T_ Natarajan, and Kamisetty R Rao. 1974. Discrete cosine transform. TC (1974).
- AI (2023a) Inflection AI. 2023a. Our policy on frontier safety. (2023). https://inflection.ai/frontier-safety
- AI (2023b) Mosaic AI. 2023b. Introducing MPT-7B: A New Standard for Open-Source, Commercially Usable LLMs. (2023). https://www.databricks.com/blog/mpt-7b
- AI (2024) Meta AI. 2024. Llama 2 - acceptable use policy. (2024). https://ai.meta.com/llama/use-policy/
- AI (2023c) Zhipu AI. 2023c. Chatglm. (2023). https://chatglm.cn/main/detail
- Akhtar and Mian (2018) Naveed Akhtar and Ajmal Mian. 2018. Threat of adversarial attacks on deep learning in computer vision: A survey. Ieee Access (2018).
- Akhtar et al. (2021) Naveed Akhtar, Ajmal Mian, Navid Kardan, and Mubarak Shah. 2021. Advances in adversarial attacks and defenses in computer vision: A survey. IEEE Access (2021).
- Alayrac et al. (2022) Jean-Baptiste Alayrac, Jeff Donahue, Pauline Luc, Antoine Miech, Iain Barr, Yana Hasson, Karel Lenc, Arthur Mensch, Katherine Millican, Malcolm Reynolds, et al. 2022. Flamingo: a visual language model for few-shot learning. NIPS (2022).
- Anthropic (2024a) Anthropic. 2024a. Claude. (2024). https://www.anthropic.com/claude
- Anthropic (2024b) Anthropic. 2024b. Claude usage policies. (2024). https://www.anthropic.com/legal/aup
@@
- Attack and Competition (2017) NIPS 2017 Adversarial Attack and Defense Competition. 2017. ImageNet-Compatible. (2017). https://www.kaggle.com/datasets/google-brain/nips-2017-adversarial-learning-development-set/data
@@
- Awadalla et al. (2023) Anas Awadalla, Irena Gao, Josh Gardner, Jack Hessel, Yusuf Hanafy, Wanrong Zhu, Kalyani Marathe, Yonatan Bitton, Samir Gadre, Shiori Sagawa, et al. 2023. Openflamingo: An open-source framework for training large autoregressive vision-language models. arXiv preprint arXiv:2308.01390 (2023).
@@
- Badjie et al. (2024) Bakary Badjie, José Cecílio, and Antonio Casimiro. 2024. Adversarial attacks and countermeasures on image classification-based deep learning models in autonomous driving systems: A systematic review. Comput. Surveys (2024).
@@
- Bagdasaryan et al. (2023) Eugene Bagdasaryan, Tsung-Yin Hsieh, Ben Nassi, and Vitaly Shmatikov. 2023. (Ab) using Images and Sounds for Indirect Instruction Injection in Multi-Modal LLMs. arXiv preprint arXiv:2307.10490 (2023).
- Bai et al. (2023a) Jinze Bai, Shuai Bai, Yunfei Chu, Zeyu Cui, Kai Dang, Xiaodong Deng, Yang Fan, Wenbin Ge, Yu Han, Fei Huang, et al. 2023a. Qwen technical report. arXiv preprint arXiv:2309.16609 (2023).
@@
- Bai et al. (2023b) Jinze Bai, Shuai Bai, Shusheng Yang, Shijie Wang, Sinan Tan, Peng Wang, Junyang Lin, Chang Zhou, and Jingren Zhou. 2023b. Qwen-vl: A versatile vision-language model for understanding, localization, text reading, and beyond. (2023).
@@
- Bai et al. (2022) Yuntao Bai, Andy Jones, Kamal Ndousse, Amanda Askell, Anna Chen, Nova DasSarma, Dawn Drain, Stanislav Fort, Deep Ganguli, Tom Henighan, et al. 2022. Training a helpful and harmless assistant with reinforcement learning from human feedback. arXiv preprint arXiv:2204.05862 (2022).
- Baidu (2023) Baidu. 2023. Ernie bot: Baidu’s knowledge-enhanced large language model built on full ai stack technology. (2023). http://research.baidu.com/Blog/index-view?id=183
@@
- Bailey et al. (2023) Luke Bailey, Euan Ong, Stuart Russell, and Scott Emmons. 2023. Image hijacks: Adversarial images can control generative models at runtime. arXiv preprint arXiv:2309.00236 (2023).
@@
- Baluja and Fischer (2017) Shumeet Baluja and Ian Fischer. 2017. Adversarial transformation networks: Learning to generate adversarial examples. arXiv preprint arXiv:1703.09387 (2017).
@@
- Bao et al. (2023) Fan Bao, Shen Nie, Kaiwen Xue, Chongxuan Li, Shi Pu, Yaole Wang, Gang Yue, Yue Cao, Hang Su, and Jun Zhu. 2023. One transformer fits all distributions in multi-modal diffusion at scale. In ICML.
@@
- Bao et al. (2022) Hangbo Bao, Wenhui Wang, Li Dong, Qiang Liu, Owais Khan Mohammed, Kriti Aggarwal, Subhojit Som, Songhao Piao, and Furu Wei. 2022. Vlmo: Unified vision-language pre-training with mixture-of-modality-experts. NIPS (2022).
@@
- Biggio et al. (2013) Battista Biggio, Igino Corona, Davide Maiorca, Blaine Nelson, Nedim Šrndić, Pavel Laskov, Giorgio Giacinto, and Fabio Roli. 2013. Evasion attacks against machine learning at test time. In ECML PKDD.
@@
- Bode et al. (2021) Lisa Bode, Dominic Lees, and Dan Golding. 2021. The digital face and deepfakes on screen.
@@
- Brendel et al. (2017) Wieland Brendel, Jonas Rauber, and Matthias Bethge. 2017. Decision-based adversarial attacks: Reliable attacks against black-box machine learning models. arXiv preprint arXiv:1712.04248 (2017).
@@
- Brock et al. (2021) Andy Brock, Soham De, Samuel L Smith, and Karen Simonyan. 2021. High-performance large-scale image recognition without normalization. In ICML.
@@
- Brown et al. (2018) Tom B Brown, Nicholas Carlini, Chiyuan Zhang, Catherine Olsson, Paul Christiano, and Ian Goodfellow. 2018. Unrestricted adversarial examples. arXiv preprint arXiv:1809.08352 (2018).
@@
- Brown et al. (2017) Tom B Brown, Dandelion Mané, Aurko Roy, Martín Abadi, and Justin Gilmer. 2017. Adversarial patch. arXiv preprint arXiv:1712.09665 (2017).
@@
- Carlini et al. (2024) Nicholas Carlini, Milad Nasr, Christopher A Choquette-Choo, Matthew Jagielski, Irena Gao, Pang Wei W Koh, Daphne Ippolito, Florian Tramer, and Ludwig Schmidt. 2024. Are aligned neural networks adversarially aligned? NIPS (2024).
@@
- Caselli et al. (2020) Tommaso Caselli, Valerio Basile, Jelena Mitrović, and Michael Granitzer. 2020. Hatebert: Retraining bert for abusive language detection in english. arXiv preprint arXiv:2010.12472 (2020).
- Chao et al. (2024) Patrick Chao, Edoardo Debenedetti, Alexander Robey, Maksym Andriushchenko, Francesco Croce, Vikash Sehwag, Edgar Dobriban, Nicolas Flammarion, George J Pappas, Florian Tramer, et al. 2024. Jailbreakbench: An open robustness benchmark for jailbreaking large language models. arXiv preprint arXiv:2404.01318 (2024).
@@
- Chao et al. (2023) Patrick Chao, Alexander Robey, Edgar Dobriban, Hamed Hassani, George J Pappas, and Eric Wong. 2023. Jailbreaking black box large language models in twenty queries. arXiv preprint arXiv:2310.08419 (2023).
@@
- Chen et al. (2022) Huanran Chen, Shitong Shao, Ziyi Wang, Zirui Shang, Jin Chen, Xiaofeng Ji, and Xinxiao Wu. 2022. Bootstrap generalization ability from loss landscape perspective. In ECCV.
@@
- Chen et al. (2023c) Huanran Chen, Yichi Zhang, Yinpeng Dong, Xiao Yang, Hang Su, and Jun Zhu. 2023c. Rethinking model ensemble in transfer-based adversarial attacks. arXiv preprint arXiv:2303.09105 (2023).
@@
- Chen et al. (2023d) Jun Chen, Deyao Zhu, Xiaoqian Shen, Xiang Li, Zechun Liu, Pengchuan Zhang, Raghuraman Krishnamoorthi, Vikas Chandra, Yunyang Xiong, and Mohamed Elhoseiny. 2023d. Minigpt-v2: large language model as a unified interface for vision-language multi-task learning. arXiv preprint arXiv:2310.09478 (2023).
@@
- Chen et al. (2017) Pin-Yu Chen, Huan Zhang, Yash Sharma, Jinfeng Yi, and Cho-Jui Hsieh. 2017. Zoo: Zeroth order optimization based black-box attacks to deep neural networks without training substitute models. In AISec.
@@
- Chen et al. (2024a) Shuo Chen, Zhen Han, Bailan He, Zifeng Ding, Wenqian Yu, Philip Torr, Volker Tresp, and Jindong Gu. 2024a. Red Teaming GPT-4V: Are GPT-4V Safe Against Uni/Multi-Modal Jailbreak Attacks? arXiv preprint arXiv:2404.03411 (2024).
@@
- Chen et al. (2019) Shang-Tse Chen, Cory Cornelius, Jason Martin, and Duen Horng Chau. 2019. Shapeshifter: Robust physical adversarial attack on faster r-cnn object detector. In ECML PKDD.
@@
- Chen et al. (2015) Xinlei Chen, Hao Fang, Tsung-Yi Lin, Ramakrishna Vedantam, Saurabh Gupta, Piotr Dollár, and C Lawrence Zitnick. 2015. Microsoft coco captions: Data collection and evaluation server. arXiv preprint arXiv:1504.00325 (2015).
@@
- Chen et al. (2023a) Xinquan Chen, Xitong Gao, Juanjuan Zhao, Kejiang Ye, and Cheng-Zhong Xu. 2023a. Advdiffuser: Natural adversarial example synthesis with diffusion models. In ICCV.
- Chen et al. (2023b) Yang Chen, Ethan Mendes, Sauvik Das, Wei Xu, and Alan Ritter. 2023b. Can language models be instructed to protect personal information? arXiv preprint arXiv:2310.02224 (2023).

### v2

@@ 7.5. Adversarial Defenses in LVLM
[(Li et al., 2024e; Zhang et al., 2024b; Chen et al., 2024a) , color=brightlavender, fill=brightlavender!15, text width=1.7cm, text=black]
@@
[(Ji et al., 2024b; Zhang et al., 2024b) , color=brightlavender, fill=brightlavender!15, text width=1.7cm, text=black]
@@
[(Zhang et al., 2023c; Li et al., 2024c) , color=brightlavender, fill=brightlavender!15, text width=1.7cm, text=black]
@@
[(Yin et al., 2023; Qi et al., 2024; Wang et al., 2024c; Gou et al., 2024) , color=brightlavender, fill=brightlavender!15, text width=1.7cm, text=black]
@@
[(Chen et al., 2023c; Phute et al., 2023; OpenAI, 2023c) , color=brightlavender, fill=brightlavender!15, text width=1.7cm, text=black]
@@
Completely filtering biased or harmful content from training corpora is infeasible, so various defenses have been developed—much like how laws guide ethical behavior when people inevitably encounter harmful information. Fig. 12 illustrates the classification of adversarial defense methods in LVLM. Defense methods during the training phase involve fine-tuning the model or inputs to enhance robustness, while those during the inference phase focus on data detection and modification. As mentioned in JBA (Qi et al., 2024) and SSA-CWA (Dong et al., 2023a), FFT (Full Fine-Tuning) through adversarial training is computationally expensive. Therefore, the cost-effective methods like fine-tuning partial parameters with LoRA (Ji et al., 2024b; Zhang et al., 2024b), prompt tuning of the inputs (Zhang et al., 2023c; Li et al., 2024c), and the inference-phase strategies are more prevalent.
@@ 7.5.1. Training Phase
Inference-phase defense can be achieved by fine-tuning models or inputs. RTVLM (Li et al., 2024e), SPA-VL (Zhang et al., 2024b), and Dress (Chen et al., 2024a) apply adversarial FFT on LVLMs with SFT (Supervised Fine-Tuning), RLHF, and NLF-based CLF (Natural Language Feedback based Conditional Reinforcement Learning), respectively. To reduce computational costs, AdvLoRA (Ji et al., 2024b) and SPA-VL (Zhang et al., 2024b) explore LoRA-based adversarial training strategies. Inspired by prompt tuning, AdvPT (Zhang et al., 2023c) and APT (Li et al., 2024c) attempt to fine-tune the context of labels in image captioning under CLIP to improve robustness.
@@ 7.5.2. Inference Phase
Inference-phase methods can be further divided into pre-processing and post-processing, based on the time of LVLM’s replying. In pre-processing, JailGuard (Zhang et al., 2023d) detects whether the input is harmful by observing the LVLM’s response to different input variations (e.g., word changes or image transformations). ESCO (Gou et al., 2024) concatenates the LVLM’s self-generated image description to the original prompt to form a new input, avoiding adversarial information in the image. SASP (Yin et al., 2023) and AdaShield (Wang et al., 2024c) insert elaborate defense prompts into user prompts, guiding the model to generate safe content. JBA (Qi et al., 2024) finds that diffusion models (Nie et al., 2022) can effectively purify jailbreak images. In post-processing, LSD (Phute et al., 2023), CM-4 (OpenAI, 2023c), and PPI (Chen et al., 2023c) use additional LLMs or the victim moddel itself to detect whether the response is harmful.
@@ 8.1.1. Traditional Adversarial Attacks
Transferability. Transfer rates of untargeted attacks have reached around 90% (even with defenses) (Chen et al., 2023d), while targeted transfer rates lag significantly (Qin et al., 2022). Targeted attacks are more demanding, and boosting their transferability is challenging. Existing transfer attacks often rely on aggregation to enhance AEs’ generalization, but extensive aggregation incurs high computational costs. Developing lightweight transfer strategies represents a meaningful research direction.
Stealthiness. Traditional -norm-based constraints often produce noise-like perturbations, which can be easily detected by attentive users (section 5.3). Therefore, designing semantically meaningful perturbations to evade from perception of humans is highly significant. Existing methods can achieve this through perceptual distance constraints, generative models, and camouflage (or watermarking). Methods with more efficiency while maintaining ASR is worth exploring.
Physical Robustness. For physical robustness, attacks mainly focuses on target function design and transformation aggregation ways (section 5.2), which seems overly monotonous. Since physically robust AEs threaten many applications, discovering new ways to enhance this robustness offers valuable insights for defenses. Moreover, if these samples also transfer across models, they may pose an even greater threat to real-world systems. In addition, physically robust AEs often require visible perturbations, while increasing naturalness is a hot spot for improving stealthiness. Is there other way to combine both—creating seemingly natural but physically threatening AEs without relying on camouflage?
Applications and Extensions. In application and extension, the intersection between AEs and interpretable machine learning is particularly significant. On one hand, the generation mechanisms of AEs provide new insights into the black-box nature of machine learning systems. On the other hand, AEs aligned with model interpretation mechanisms pose greater threats across various social domains. Furthermore, machine learning applications ranging from pioneering techniques like style transfer (Shan et al., 2023) and face swapping (Dong et al., 2023b) to current popular multimodal generation systems (e.g., text-to-image models) all face potential adversarial attack risks. Investigating the differences and connections among AEs across different tasks—especially multimodal tasks—represents a meaningful research direction.
@@ 8.1.2. Adversarial Attacks on LVLM
Adversariality. In LVLM scenarios, attacks on textual modalities remain understudied compared to visual modalities. Strengthening text-based attacks and exploring the vulnerability in cross-modality links could be promising directions. As all modalities contain information redundancy, precisely utilizing multimodal features to identify effective perturbation strategies warrants investigation. Beyond conventional adversarial perturbations and prompt manipulation, developing novel attack surfaces (e.g., CoT attacks (Wang et al., 2024d; Kuo et al., 2025) and resource-exhaustion attacks (Gao et al., 2024b)) holds significant research value.
Transferability. For LVLMs, both the visual encoder and LLM predominantly adopt the Transformer architecture, which may potentially facilitate transfer attacks to some extent. However, enhancing cross-model transferability of AEs remains a fundamental challenge in LVLM scenarios. Furthermore, as LVLMs vary in parameter scales, AEs transferable across differently scaled models would pose significant threats. Similar to traditional transfer attacks, LVLM-based attacks often employ aggregation to improve generalization. Given the high parameter count of LVLMs, aggregation becomes considerably challenging, making lightweight generation methods particularly crucial in such contexts.
Physical Robustness. Current adversarial attacks targeting LVLMs in physical environments remain underexplored. As LVLMs become increasingly integrated into real-world applications like autonomous driving, they may trigger renewed security concerns. Generating physically robust AEs in LVLM scenarios is a topic of practical significance.
Generation Efficiency. The massive parameter sizes and paid commercial APIs of LVLMs impose high costs on attackers. How to cost-effectively generate AEs remains a key challenge. Besides, in real-time LVLM applications like autonomous driving and pedestrian recognition, enhancing generation speed emerges as a critical requirement.
Applications and Extensions. LVLMs are gradually integrating into various aspects of daily life. Existing methods have demonstrated attacks on LVLM-based shopping agents (Wu et al., 2024) and applications of AEs for copyright protection (Shan et al., 2023). Exploring AEs on different LVLM downstream applications can provide valuable insights for their deployment.
Benchmarks. Accurately and efficiently assessing the robustness of LVLMs is a serious challenge. While numerous safety related benchmarks (Shen et al., 2024; Liu et al., 2023d; Li et al., 2024a; Luo et al., 2024b; Zhang et al., 2024a; Li et al., 2024e; Tu et al., 2023; Sun et al., 2023b) for LVLMs have been proposed recently, there is still a lack of a widely accepted, unified, and comprehensive evaluation system.
@@ 8.2. Conclusion
This article reviews the development of visual adversarial attacks over the past decade. As Qi et al. (Qi et al., 2024) note, adversarial attacks are shifting from classification-focused approaches to broader applications across LLMs. The article is thus divided into two parts: 1) traditional adversarial attacks, and 2) LVLM adversarial attacks.
The first part summarizes adversariality, transferability, and generalization, explaining the causes of adversariality and transferability, the roles of AEs, traits of transferability, and types of generalization. It then defines the problem and introduces threat models, victim models, datasets, and evaluation metrics, classifying traditional attacks into two phases: basic strategies, which explore various attack paradigms, and attack enhancement, aimed at boosting effectiveness. Motivations of the second phase fall into four types: improving transferability, physical robustness, stealthiness, and generation efficiency. This part concludes with an overview of the application of attacks across different tasks.
@@
Finally, the article discusses future research directions, including adversariality, transferability, physical robustness, stealth, generation efficiency, and applications, aiming to provide insights for future works in visual adversarial attacks.
@@ 9. Acknowledgments
This research was supported by the National Key R&D Program of China (No. 2022YFB2702000), the Natural Science Foundation of Jiangsu Province, China (No. BK20220075), the National Natural Science Foundation of China (No. 62132008, U22B2030, 62472218). We thank the Collaborative Innovation Center of Novel Software Technology and Industrialization for their support. We also thank Xiaogang Xu from The Chinese University of Hong Kong, the project leader, for his contribution to this paper.
@@
- 3) (2017) NIPS 2017 Defense Competition (Rank 3). 2017. NIPS-r3. https://github.com/anlthms/nips-2017/tree/master/mmd
- Abid et al. (2021) Abubakar Abid, Maheen Farooqi, et al. 2021. Large language models associate Muslims with violence. Nature Machine Intelligence.
- Agnihotri et al. (2024) Shashank Agnihotri, Steffen Jung, et al. 2024. CosPGD: an efficient white-box adversarial attack for pixel-wise prediction tasks. In ICML.
- Agrawal et al. (2019) Harsh Agrawal, Karan Desai, Yufei Wang, Xinlei Chen, Rishabh Jain, et al. 2019. Nocaps: Novel object captioning at scale. In ICCV.
- Ahmed et al. (1974) Nasir Ahmed, T Natarajan, and Kamisetty R Rao. 1974. Discrete Cosine Transfom. IEEE Trans. Comput.
- AI (2023a) Inflection AI. 2023a. Our policy on frontier safety. https://inflection.ai/frontier-safety
- AI (2024) Meta AI. 2024. Llama 2 - acceptable use policy. https://ai.meta.com/llama/use-policy/
- AI (2023b) Zhipu AI. 2023b. Chatglm. https://chatglm.cn/main/detail
- Akhtar and Mian (2018) Naveed Akhtar and Ajmal Mian. 2018. Threat of adversarial attacks on deep learning in computer vision: A survey. IEEE Access.
- Akhtar et al. (2021) Naveed Akhtar, Ajmal Mian, et al. 2021. Advances in adversarial attacks and defenses in computer vision: A survey. IEEE Access.
- Alayrac et al. (2022) Jean-Baptiste Alayrac, Jeff Donahue, Pauline Luc, et al. 2022. Flamingo: a visual language model for few-shot learning. In NIPS.
- Anthropic (2024a) Anthropic. 2024a. Claude. https://www.anthropic.com/claude
- Anthropic (2024b) Anthropic. 2024b. Claude usage policies. https://www.anthropic.com/legal/aup
@@
- Attack and Competition (2017) NIPS 2017 Adversarial Attack and Defense Competition. 2017. ImageNet-Compatible. https://www.kaggle.com/datasets/google-brain/nips-2017-adversarial-learning-development-set/data
@@
- Awadalla et al. (2023) Anas Awadalla et al. 2023. Openflamingo: An open-source framework for training large autoregressive vision-language models. arXiv:2308.01390.
@@
- Badjie et al. (2024) Bakary Badjie, José Cecílio, and Antonio Casimiro. 2024. Adversarial attacks and countermeasures on image classification-based deep learning models in autonomous driving systems: A systematic review. Comput. Surveys.
@@
- Bagdasaryan et al. (2023) Eugene Bagdasaryan et al. 2023. (Ab) using Images and Sounds for Indirect Instruction Injection in Multi-Modal LLMs. arXiv:2307.10490.
- Bahramali et al. (2021) Alireza Bahramali, Milad Nasr, et al. 2021. Robust adversarial attacks against DNN-based wireless communication systems. In ACM SIGSAC.
@@
- Bai et al. (2023b) Jinze Bai et al. 2023b. Qwen-vl: A versatile vision-language model for understanding, localization, text reading, and beyond. arXiv:2308.12966.
- Bai et al. (2023a) Jinze Bai, Shuai Bai, Yunfei Chu, Zeyu Cui, Kai Dang, Xiaodong Deng, et al. 2023a. Qwen technical report. arXiv:2309.16609.
@@
- Bai et al. (2022) Yuntao Bai et al. 2022. Training a helpful and harmless assistant with reinforcement learning from human feedback. arXiv:2204.05862.
- Baidu (2023) Baidu. 2023. Ernie bot Introduction Report. http://research.baidu.com/Blog/index-view?id=183
@@
- Bailey et al. (2023) Luke Bailey, Euan Ong, et al. 2023. Image hijacks: Adversarial images can control generative models at runtime. arXiv:2309.00236.
@@
- Baluja and Fischer (2017) Shumeet Baluja and Ian Fischer. 2017. Adversarial transformation networks: Learning to generate adversarial examples. arXiv:1703.09387.
@@
- Bao et al. (2023) Fan Bao, Shen Nie, Kaiwen Xue, Chongxuan Li, et al. 2023. One transformer fits all distributions in multi-modal diffusion at scale. In ICML.
@@
- Bao et al. (2022) Hangbo Bao, Wenhui Wang, Li Dong, et al. 2022. Vlmo: Unified vision-language pre-training with mixture-of-modality-experts. In NIPS.
@@
- Biggio et al. (2013) Battista Biggio, Igino Corona, Davide Maiorca, et al. 2013. Evasion attacks against machine learning at test time. In ECML PKDD.
- Blum et al. (2022) Avrim Blum, Omar Montasser, et al. 2022. Boosting barely robust learners: A new perspective on adversarial robustness. In NIPS.
@@
- Bode et al. (2021) Lisa Bode, Dominic Lees, and Dan Golding. 2021. The digital face and deepfakes on screen. Convergence.
@@
- Brendel et al. (2017) W Brendel et al. 2017. Decision-based adversarial attacks: Reliable attacks against black-box machine learning models. arXiv:1712.04248.
@@
- Brock et al. (2021) Andy Brock, Soham De, Samuel L Smith, et al. 2021. High-performance large-scale image recognition without normalization. In ICML.
@@
- Brown et al. (2018) Tom B Brown, Nicholas Carlini, Chiyuan Zhang, Catherine Olsson, et al. 2018. Unrestricted adversarial examples. arXiv:1809.08352.
@@
- Brown et al. (2017) Tom B Brown, Dandelion Mané, Aurko Roy, Martín Abadi, and Justin Gilmer. 2017. Adversarial patch. arXiv:1712.09665.
@@
- Carlini et al. (2024) Nicholas Carlini, Milad Nasr, Christopher A Choquette-Choo, et al. 2024. Are aligned neural networks adversarially aligned?. In NIPS.
@@
- Caselli et al. (2020) Tommaso Caselli et al. 2020. Hatebert: Retraining bert for abusive language detection in english. arXiv:2010.12472.
- Chao et al. (2024) Patrick Chao et al. 2024. Jailbreakbench: An open robustness benchmark for jailbreaking large language models. arXiv:2404.01318.
@@
- Chao et al. (2023) Patrick Chao, Alexander Robey, et al. 2023. Jailbreaking black box large language models in twenty queries. arXiv:2310.08419.
@@
- Chen et al. (2022) Huanran Chen, Shitong Shao, Ziyi Wang, Zirui Shang, et al. 2022. Bootstrap generalization ability from loss landscape perspective. In ECCV.
@@
- Chen et al. (2023d) Huanran Chen, Yichi Zhang, et al. 2023d. Rethinking model ensemble in transfer-based adversarial attacks. arXiv:2303.09105.
@@
- Chen et al. (2023a) Jun Chen et al. 2023a. Minigpt-v2: large language model as a unified interface for vision-language multi-task learning. arXiv:2310.09478.
@@
- Chen et al. (2017) P Chen et al. 2017. Zoo: Zeroth order optimization based black-box attacks to deep neural networks without training substitute models. In AISec.
@@
- Chen et al. (2024b) Shuo Chen, Zhen Han, et al. 2024b. Red Teaming GPT-4V: Are GPT-4V Safe Against Uni/Multi-Modal Jailbreak Attacks? arXiv:2404.03411.
@@
- Chen et al. (2019) Shang-Tse Chen, Cory Cornelius, et al. 2019. Shapeshifter: Robust physical adversarial attack on faster r-cnn object detector. In ECML PKDD.
@@
- Chen et al. (2015) Xinlei Chen, Hao Fang, et al. 2015. Microsoft coco captions: Data collection and evaluation server. arXiv:1504.00325.
@@
- Chen et al. (2023b) Xinquan Chen, Xitong Gao, Juanjuan Zhao, et al. 2023b. Advdiffuser: Natural adversarial example synthesis with diffusion models. In ICCV.